Can someone help me solve this problem?
I need to find two solutions for the following equation:
When x number of units are sold, the price of each unit (in dollars) is given by p = -x/2 + 75. I need to find the unit price when the following quantities are sold: 9 and 11.
Thanks,
Sher

Answer by stanbon(29603)

p = -x/2 + 75. I need to find the unit price when the following quantities are sold: 9 and 11.
--------------
P(9)=-9/2 + 75
=-4.5+75
=71.5
------------
p(11)=-11/2 + 75
=-5.5 + 75
=69.5
Cheers,
Stan H.